OpenStudy (osanseviero):

Fast question: x=1/(2y-3) What is y?

OpenStudy (osanseviero):

\[\frac{1 }{2y-3 }\]

OpenStudy (anonymous):

2y-3=1/x 2y=1/x+3 => y=(1+3x)/2x